OpenStudy (anonymous):

Simplify (3xy^2)(4xy)(2xy)^3.

OpenStudy (welshfella):

deal with the numbers firat what is 3*4*2^3 ?

OpenStudy (anonymous):

24^3

OpenStudy (anonymous):

96x^5 y^6

OpenStudy (anonymous):

24^3=13,824

OpenStudy (welshfella):

no its only the 2 which is taken to the power 3

OpenStudy (welshfella):

@princess_hannahleigh please don't just give the answer. Read the Code of Conduct

OpenStudy (anonymous):

I need help smh

OpenStudy (welshfella):

its 3*4*2*2*2 (2^3 = 2*2*2)

OpenStudy (luigi0210):

He is trying to help, by giving you general steps and instruction. Giving direct answers isn't considered helping though.

OpenStudy (anonymous):

2^3 is 8

OpenStudy (anonymous):